Home‎ > ‎Crew 136‎ > ‎

225-CommanderReport

Commander Report
02/25/2014
Corentin Liber

We had the visit of a Russian media crew today.  They followed us during our usual activities and
left around 4pm.

In the morning, we begun with an EVA for our balloon project and finally launched it.  We got some
footage from the GoPro fixed to it and a lot of data from different captors.  It will now be
analyzed during the next days.

In the afternoon, the second EVA team went to collect some additional data in the streambed we're
studying because of some inconsistence with the one previously recorded.  A corrected model will be
implemented tomorrow.  The geologist also finished his sampling in the area nearby and is ready for
the next zone tomorrow.

We'll probably go out to observe the sky tonight if the weather conditions allow it.

-------------------------------

CommanderCheckIn Report
02/25/2014
Thibaud Herriau


Crew Physical Health:
All are well.

Eva Depart Time:
10am
Eva Return Time:
11:30am
Narrative Of Field Mission Results:
Launch of the balloon.
Retrieving of the experiment's data.

Eva Depart Time:
2:30pm
Eva Return Time:
4pm
Narrative Of Field Mission Results:
Streambed measurements.
Mineral sampling in a new area.

Hab Maint:
No more garbage bags. Crew137 should bring new ones.
Rep Sched:
None
Tomorrow Plan:
Long range EVA with ATV in the afternoon to explore a new area for mineral
sampling.
Treatment of the fresh data of the balloon experiment.
Modeling of the streambed.
Inventory:
None
Support Req:
None
Miscelleneous:
None